A buffer is prepared by adding 0.15 mol of lactic acid, CH3CHOHCOOH, and 0.20 mol of sodium lactate, Na+CH3CHOHCOO - , to sufficient water to make 1.0 L of buffer solution. The Ka of lactic acid is 1.4 x 10 - 4. Calculate the pH of the buffer.
Answer #1

Ka = 1.4 x 10^-4

pKa = -log Ka

pKa = -log (1.4 x 10^-4 )

pKa = 3.85

pH = pKa + log [sodium lactate / lactic acid ]

pH = 3.85 + log (0.20 / 0.15)

pH = 3.98